# Correlated Noise in Epoch-Based Stochastic Gradient Descent: Implications for Weight Variances

### v2

@@ C.3 Approximation of the exact formula
It is straightforward to show that the magnitude of the eigenvalues of $\vec{D}$ is strictly smaller than one, $|\Lambda_{\pm}|<1$, under the conditions $0<\eta\lambda<2(1+\beta)$ and $10\leq\beta<$ 1.
@@
Inserting the eigendecomposition of $\bm{\mathbf{D}}$ into the expression $\vec{D}^{M}\vec{e_{1}}$ yields
@@
where $\left\lVert\cdot\right\rVert_{\infty}$ is denoting the maximum norm $\left\lVert\bm{\mathbf{x}}\right\rVert_{\infty}\coloneqq\max_{i}|x_{i}|$ for a vector $\vec{x}$ or its induced matrix norm $\left\lVert\bm{\mathbf{A}}\right\rVert_{\infty}\coloneqq\max_{i}\sum_{j}|a_{ij}|$ for a matrix $\vec{A}$.
@@
$\left\lVert(\bm{\mathbf{1}}-\vec{D})^{-1}\right\rVert_{\infty}\leq\frac{4}{\eta\lambda}$ (70)
under the assumption that $10\leq\beta<$ 1 and $0<\eta\lambda<2(1+\beta)$. From here it is straightforward to show that
@@
where $\tilde{c}$ is a factor of order unity under the constraints $10\leq\beta<$ 1 and $0<\eta\lambda<2(1+\beta)$. By substituting this result back into Equation 55 one directly sees that a comparison to the approximation yields
@@
To obtain the relations for small hessian eigenvalues, we perform a Taylor expansion with respect to $\lambda$ with the help of computer algebra. We neglect the terms which are at least of order lambda. Numerical study indicates that these relations hold under the mentioned assumption of $\lambda\lesssim\lambda_{\textrm{cross}}$ and $M(1-\beta)\gg 1$.
@@
C.4 Satisfying the assumptions of the correlation time relation
In this section we want to show that the weight and velocity variances resulting from stochastic gradient descent as described above and in Section 4.3 satisfies the necessary assumptions (i) to (iii) of Section 4.2 such that the velocity correlation time is equal to $\tau_{i}=2\sigma_{\theta,i}^{2}/\sigma_{v,i}^{2}$. Validity of assumption (i) existence and finiteness of $\bm{\mathbf{\Sigma}}\coloneqq$ cov$(\bm{\mathbf{\theta}}_{k},\bm{\mathbf{\theta}}_{k})$, $\bm{\mathbf{\Sigma_{v}}}\coloneqq$ cov$(\bm{\mathbf{v}}_{k},\bm{\mathbf{v}}_{k})$, and $\langle\bm{\mathbf{\theta}}\rangle$ can be inferred from the calculation presented in Section C.1. Therefore, we concentrate on assumption (ii) $\lim_{n\to\infty}\textrm{cov}\bigr{(}\bm{\mathbf{\theta}}_{k},\bm{\mathbf{\theta}}_{k+n}\bigr{)}=0$ and (iii) $\lim_{n\to\infty}n\cdot\textrm{cov}\bigr{(}\bm{\mathbf{\theta}}_{k},\bm{\mathbf{\theta}}_{k+n}-\bm{\mathbf{\theta}}_{k+n+1}\bigr{)}=0$. We consider the one dimensional case, but the extension to the multidimensional case is straightforward. Additionally $\langle\vec{\theta}\rangle=0$ (see Section C.1) and, therefore, the remaining two assumptions (ii) and (iii) can be written as $\lim_{m\rightarrow\infty}{\langle\theta_{k}\theta_{k+m}\rangle}=0$ and $\lim_{m\rightarrow\infty}{m\bigl{(}\langle\theta_{k}\theta_{k+m}\rangle-\langle\theta_{k}\theta_{k+m+1}\rangle\bigr{)}}=0$.
We will now show that for stochastic gradient descent under the assumptions of Section 4.3 the more restrictive relation $\lim_{m\rightarrow\infty}{m\langle\theta_{k}\theta_{k+m}\rangle}=0$ is satisfied, from which follows (ii) and (iii). Following Section C.2 and using the same notation, we have the relation
$\langle\vec{x}_{k}\vec{x}_{k-m}^{\top}\rangle=\bm{\mathbf{D}}\langle\vec{x}_{k-1}\vec{x}_{k-m}^{\top}\rangle-\eta\vec{e_{1}}\langle\delta g_{k}\vec{x}_{k-m}^{\top}\rangle\ ,$ (76)
where $\vec{x}_{k}\coloneqq\bigl{(}\theta_{k}\quad\theta_{k-1}\bigr{)}^{\top}$. For $m>M$, with $M$ being the number of batches per epoch, the correlation with the noise term on the right hand side of Equation 76 is equal to zero as discussed in Section C.2. By iterating Equation 76, for $m>M$ we have
$\langle\vec{x}_{k}\vec{x}_{k-m}^{\top}\rangle$ $=\bm{\mathbf{D}}^{m-M-1}\langle\vec{x}_{k-m+M+1}\vec{x}_{k-m}^{\top}\rangle$
$=\bm{\mathbf{D}}^{m-M-1}\langle\vec{x}_{k}\vec{x}_{k-M-1}^{\top}\rangle\ .$ (77)
As described in Section C.2, the magnitude of both eigenvalues of $\vec{D}$ is strictly smaller than one. This implies that there exists a matrix norm $\left\lVert\cdot\right\rVert_{\vec{D}}$ such that $\left\lVert\vec{D}\right\rVert_{\vec{D}}<1$ from which one can deduce
$\left\lVert\langle\vec{x}_{k}\vec{x}_{k-m}^{\top}\rangle\right\rVert_{\vec{D}}\leq\left\lVert\bm{\mathbf{D}}\right\rVert_{\vec{D}}^{m-M-1}\cdot\left\lVert\langle\vec{x}_{k}\vec{x}_{k-M-1}^{\top}\rangle\right\rVert_{\vec{D}}\ .$ (78)
@@ C.4 Satisfying the assumptions of the correlation time relation
$\lim_{m\to\infty}m\left\lVert\langle\vec{x}_{k}\vec{x}_{k-m}^{\top}\rangle\right\rVert_{\vec{D}}$ $\leq\textrm{const}\cdot\lim_{m\to\infty}m\left\lVert\bm{\mathbf{D}}\right\rVert_{\vec{D}}^{m-M-1}$
@@
Appendix D Calculation of the Correlation Time Relation
We want to prove the relation
$\frac{2\sigma_{\theta,i}^{2}}{\sigma_{v,i}^{2}}=\frac{\sum_{n=1}^{\infty}n\,\langle v_{k,i}v_{k+n,i}\rangle}{\sum_{n=1}^{\infty}\langle v_{k,i}v_{k+n,i}\rangle}\ .$ (82)
We assume that that $\langle\vec{\theta}\rangle$, $\bm{\mathbf{\Sigma_{\theta}}}$ and $\bm{\mathbf{\Sigma_{v}}}$ exist and are finite. Without loss of generality, let $\langle\vec{\theta}\rangle=0$. We consider only the one-dimensional case. For the multidimensional case, simply apply the proof in the direction of any basis vector individually. Note, that the relation still holds if $[\bm{\mathbf{\Sigma}},\bm{\mathbf{\Sigma_{v}}}]\neq 0$. In this case, $\sigma_{\theta,i}^{2}$ and $\sigma_{v,i}^{2}$ would just be the variances of the weight and the velocity in the given direction but no longer necessarily eigenvalues of $\bm{\mathbf{\Sigma}}$ and $\bm{\mathbf{\Sigma_{v}}}$.
The remaining two assumptions (ii) and (iii) of Section 4.2 can now be written as
@@ Appendix D Calculation of the Correlation Time Relation
Therefore, we can rewrite Equation 85 as follows
@@
where we first shifted the index within the expectation value and then restructured the sum by defining $m\coloneqq\max(i,j)-1$ and $n\coloneqq j-i$. We now take the limit of $J\rightarrow\infty$ and because of Equation 83 and the assumption of a finite $\sigma_{\theta}^{2}$ we have
$\sum_{m=0}^{\infty}\sum_{n=-m}^{m}\langle v_{k}v_{k+n}\rangle$ <$\infty$< ∞ (88)
@@
We note that $\langle v_{k}v_{k+n}\rangle=\langle v_{k}v_{k-n}\rangle$ because we can shift the index, and the two factors commute. Substituting this relation into Equation 89 yields
@@
where nearly all terms cancel each other again due to the fact that we can shift the index within the expectation value. By taking the limit $m\rightarrow\infty$ and using the assumptions (ii) and (iii) (Equations 83 and 84) we have
@@
Finally, by dividing Equation 92 by Equation 90 we arrive at the final expression
@@
Appendix E Calculation of the Noise Autocorrelation
We want to calculate the autocorrelation function of epoch-based SGD for a fixed weight vector $\bm{\mathbf{\theta}}$ and under the assumption that the total number of examples is an integer multiple of the number of examples per batch. For that we repeat the following definitions:
$\bm{\mathbf{\delta g}}_{k}(\bm{\mathbf{\theta}})$ $\coloneqq\frac{1}{S}\sum_{n\in\mathcal{B}_{k}}\bm{\mathbf{\nabla}}\bigl{(}l(\bm{\mathbf{\theta}},x_{n})-L(\bm{\mathbf{\theta}})\bigr{)}$ (94)
@@ Appendix E Calculation of the Noise Autocorrelation
[…]
@@
$\bm{\mathbf{\delta g}}_{k}(\bm{\mathbf{\theta}})$ $=\frac{1}{S}\sum_{n\in\mathcal{B}_{k}}\bm{\mathbf{\delta g_{e}}}(n,\bm{\mathbf{\theta}})$
$=\frac{1}{S}\sum_{n=1}^{N}\bm{\mathbf{\delta g_{e}}}(n,\bm{\mathbf{\theta}})s_{k}^{n}$ (99)
@@
=$=\begin{cases}1\ \text{ if }~{}n\in\mathcal{B}_{k}\\ 0\ \text{ if }~{}n\notin\mathcal{B}_{k}\end{cases}$ (100)
$\bm{\mathbf{\delta g_{e}}}(n,\bm{\mathbf{\theta}})$ $\coloneqq\bm{\mathbf{\nabla}}\bigl{(}l(\bm{\mathbf{\theta}},x_{n})-L(\bm{\mathbf{\theta}})\bigr{)}\ .$ (101)
@@
$\textrm{cov}\bigl{(}\vec{\delta g}_{k}(\vec{\theta}),\vec{\delta g}_{k+h}(\vec{\theta})\bigr{)}$ $=\mathbb{E}\left[\bm{\mathbf{\delta g}}_{k}(\bm{\mathbf{\theta}})\bm{\mathbf{\delta g}}_{k+h}(\bm{\mathbf{\theta}})^{\top}\right]$
$=\frac{1}{S^{2}}\sum_{n,\tilde{n}=1}^{N}\bm{\mathbf{\delta g_{e}}}(n,\bm{\mathbf{\theta}})\bm{\mathbf{\delta g_{e}}}(\tilde{n},\bm{\mathbf{\theta}})^{\top}\mathbb{E}\left[s_{k}^{n}s_{k+h}^{\tilde{n}}\right]\ .$ (102)
@@
The first term of the right hand side of Equation 105 is the probability that a given example occurs in a batch, assuming that we already know one of the examples of that batch.
$\textrm{P}\!\left(s_{k}^{\tilde{n}}=1\mid s_{k}^{n}=1\right)$ $=\textrm{P}\!\left(s_{k}^{n}=1\mid s_{k}^{n}=1\right)\cdot\,\delta_{n,\tilde{n}}+\textrm{P}\!\left(s_{k}^{\tilde{n}}=1\mid s_{k}^{n}=1,n\neq\tilde{n}\right)\,(1-\delta_{n,\tilde{n}})$
$=1\cdot\,\delta_{n,\tilde{n}}+\frac{S-1}{N-1}\,(1-\delta_{n,\tilde{n}})$
$=\frac{N-S}{N-1}\,\delta_{n,\tilde{n}}+\textrm{const.}$. (106)
The second term of Equation 105 is multiplied by $(1-\delta_{h,0})$. Therefore, we assume $h\geq 1$ for the following argument. That is, we want to know the probabilities under the assumption that we are comparing examples from different batches. If the two batches are still from the same epoch, examples cannot repeat as the total number of examples is an integer multiple of the number of examples per batch and because of that every example is shown only once per epoch. Therefore, for $h\geq 1$ holds:
$\textrm{P}\!\left(s_{k+h}^{\tilde{n}}=1\mid s_{k}^{n}=1,\textrm{ep}(k)=\textrm{ep}(k+h)\right)$ $=0\cdot\delta_{n,\tilde{n}}+\frac{S}{N-1}\,(1-\delta_{n,\tilde{n}})$
$=-\frac{S}{N-1}\delta_{n,\tilde{n}}+\textrm{const.}$. (107)
@@
$\mathbb{E}\left[s_{k}^{n}s_{k+h}^{\tilde{n}}\right]$ $=\delta_{n,\tilde{n}}\,\frac{S}{N}\frac{N-S}{N-1}\left(\delta_{h,0}-\mathbf{1}_{\{1,\dots,M\}}(h)\frac{S}{N-S}\frac{M-h}{M}\right)+\textrm{const.}$.
$=\delta_{n,\tilde{n}}\,S^{2}\left(\frac{1}{S}-\frac{1}{N}\right)\frac{}{}\frac{1}{N-1}\left(\delta_{h,0}-\mathbf{1}_{\{1,\dots,M\}}(h)\frac{M-h}{M\,(M-1)}\right)+\textrm{const.}$. (110)
@@
$\sum_{n,\tilde{n}=1}^{N}\bm{\mathbf{\delta g_{e}}}(n,\bm{\mathbf{\theta}})\bm{\mathbf{\delta g_{e}}}(\tilde{n},\bm{\mathbf{\theta}})^{\top}\,\delta_{n,\tilde{n}}$ $=\sum_{n=1}^{N}\bm{\mathbf{\delta g_{e}}}(n,\bm{\mathbf{\theta}})\bm{\mathbf{\delta g_{e}}}(n,\bm{\mathbf{\theta}})^{\top}$
$\eqqcolon(N-1)\,\bm{\mathbf{C}}_{0}(\bm{\mathbf{\theta}})\ ,$ (111)
$\sum_{n,\tilde{n}=1}^{N}\bm{\mathbf{\delta g_{e}}}(n,\bm{\mathbf{\theta}})\bm{\mathbf{\delta g_{e}}}(\tilde{n},\bm{\mathbf{\theta}})^{\top}\cdot 1$ $=\left(\sum_{n=1}^{N}\bm{\mathbf{\delta g_{e}}}(n,\bm{\mathbf{\theta}})\right)\left(\sum_{n=1}^{N}\bm{\mathbf{\delta g_{e}}}(n,\bm{\mathbf{\theta}})^{\top}\right)$
$=\bm{\mathbf{\nabla}}\bigl{(}L(\bm{\mathbf{\theta}})-L(\bm{\mathbf{\theta}})\bigr{)}\bm{\mathbf{\nabla}}^{\top}\bigl{(}L(\bm{\mathbf{\theta}})-L(\bm{\mathbf{\theta}})\bigr{)}$
@@
we can insert the expectation value $\mathbb{E}\left[s_{k}^{n}s_{k+h}^{\tilde{n}}\right]$ into Equation 102 and arrive at the final expression:

### v3

@@ B.3 Approximation of the exact formula
It is straightforward to show that the magnitude of the eigenvalues of $\vec{D}$ is strictly smaller than one, $|\Lambda_{\pm}|<1$, under the conditions $0<\eta\lambda<2(1+\beta)$ and $0\leq\beta<1$.
@@
Inserting the eigendecomposition of $\boldsymbol{\mathbf{D}}$ into the expression $\vec{D}^{M}\vec{e_{1}}$ yields
@@
where $\left\lVert\cdot\right\rVert_{\infty}$ is denoting the maximum norm $\left\lVert\boldsymbol{\mathbf{x}}\right\rVert_{\infty}\coloneqq\max_{i}|x_{i}|$ for a vector $\vec{x}$ or its induced matrix norm $\left\lVert\boldsymbol{\mathbf{A}}\right\rVert_{\infty}\coloneqq\max_{i}\sum_{j}|a_{ij}|$ for a matrix $\vec{A}$.
@@
$\left\lVert(\boldsymbol{\mathbf{1}}-\vec{D})^{-1}\right\rVert_{\infty}\leq\frac{4}{\eta\lambda}$ (64)
under the assumption that $0\leq\beta<1$ and $0<\eta\lambda<2(1+\beta)$. From here it is straightforward to show that
@@
where $\tilde{c}$ is a factor of order unity under the constraints $0\leq\beta<1$ and $0<\eta\lambda<2(1+\beta)$. By substituting this result back into Equation˜49 one directly sees that a comparison to the approximation yields
@@
To obtain the relations for small Hessian eigenvalues, we perform a Taylor expansion with respect to $\lambda$ with the help of computer algebra. We neglect the terms which are at least of order $\lambda$. Numerical study indicates that these relations hold under the mentioned assumption of $\lambda\lesssim\lambda_{\textrm{cross}}$ and $M(1-\beta)\gg 1$.
@@
B.4 Satisfying the assumptions of the correlation time relation
[…]
We will now show that for stochastic gradient descent under the assumptions of Section˜4.3 the more restrictive relation $\lim_{m\rightarrow\infty}{m\langle\theta_{k}\theta_{k+m}\rangle}=0$ is satisfied, from which follows (ii) and (iii). Following Section˜B.2 and using the same notation, we have the relation
$\langle\vec{x}_{k}\vec{x}_{k-m}^{\top}\rangle=\boldsymbol{\mathbf{D}}\langle\vec{x}_{k-1}\vec{x}_{k-m}^{\top}\rangle-\eta\vec{e_{1}}\langle\delta g_{k}\vec{x}_{k-m}^{\top}\rangle\ ,$ (70)
where $\vec{x}_{k}\coloneqq\bigl(\theta_{k}\quad\theta_{k-1}\bigr)^{\top}$. For $m>M$, with $M$ being the number of batches per epoch, the correlation with the noise term on the right hand side of Equation˜70 is equal to zero as discussed in Section˜B.2. By iterating Equation˜70, for $m>M$ we have
$\langle\vec{x}_{k}\vec{x}_{k-m}^{\top}\rangle$ $=\boldsymbol{\mathbf{D}}^{m-M-1}\langle\vec{x}_{k-m+M+1}\vec{x}_{k-m}^{\top}\rangle$
$=\boldsymbol{\mathbf{D}}^{m-M-1}\langle\vec{x}_{k}\vec{x}_{k-M-1}^{\top}\rangle\ .$ (71)
As described in Section˜B.2, the magnitude of both eigenvalues of $\vec{D}$ is strictly smaller than one. This implies that there exists a matrix norm $\left\lVert\cdot\right\rVert_{\vec{D}}$ such that $\left\lVert\vec{D}\right\rVert_{\vec{D}}<1$ from which one can deduce
$\left\lVert\langle\vec{x}_{k}\vec{x}_{k-m}^{\top}\rangle\right\rVert_{\vec{D}}\leq\left\lVert\boldsymbol{\mathbf{D}}\right\rVert_{\vec{D}}^{m-M-1}\cdot\left\lVert\langle\vec{x}_{k}\vec{x}_{k-M-1}^{\top}\rangle\right\rVert_{\vec{D}}\ .$ (72)
@@ B.4 Satisfying the assumptions of the correlation time relation
$\lim_{m\to\infty}m\left\lVert\langle\vec{x}_{k}\vec{x}_{k-m}^{\top}\rangle\right\rVert_{\vec{D}}$ $\leq\textrm{const}\cdot\lim_{m\to\infty}m\left\lVert\boldsymbol{\mathbf{D}}\right\rVert_{\vec{D}}^{m-M-1}$
@@
Appendix C Calculation of the Correlation Time Relation
We want to prove Theorem˜4.2, that is, the relation
$\frac{2\sigma_{\theta,i}^{2}}{\sigma_{v,i}^{2}}=\frac{\sum_{n=1}^{\infty}n\,\langle v_{k,i}v_{k+n,i}\rangle}{\sum_{n=1}^{\infty}\langle v_{k,i}v_{k+n,i}\rangle}\ ,$ (76)
[…]
We assume that that $\langle\vec{\theta}\rangle$, $\boldsymbol{\mathbf{\Sigma_{\theta}}}$ and $\boldsymbol{\mathbf{\Sigma_{v}}}$ exist and are finite. Without loss of generality, let $\langle\vec{\theta}\rangle=0$. We consider only the one-dimensional case. For the multidimensional case, simply apply the proof in the direction of any basis vector individually. Note, that the relation still holds if $\boldsymbol{\mathbf{\Sigma}},\boldsymbol{\mathbf{\Sigma_{v}}}]\neq 0$. In this case, $\sigma_{\theta,i}^{2}$ and $\sigma_{v,i}^{2}$ would just be the variances of the weight and the velocity in the given direction but no longer necessarily eigenvalues of $\boldsymbol{\mathbf{\Sigma}}$ and $\boldsymbol{\mathbf{\Sigma_{v}}}$.
The remaining two assumptions (ii) and (iii) of Theorem˜4.2 can now be written as
@@ Appendix C Calculation of the Correlation Time Relation
Therefore, we can rewrite Equation˜79 as follows
@@
where we first shifted the index within the expectation value and then restructured the sum by defining $m\coloneqq\max(i,j)-1$ and $n\coloneqq j-i$. We now take the limit of $J\rightarrow\infty$ and because of Equation˜77 and the assumption of a finite $\sigma_{\theta}^{2}$ we have
$\sum_{m=0}^{\infty}\sum_{n=-m}^{m}\langle v_{k}v_{k+n}\rangle$ $<\infty$ (82)
@@
We note that $\langle v_{k}v_{k+n}\rangle=\langle v_{k}v_{k-n}\rangle$ because we can shift the index, and the two factors commute. Substituting this relation into Equation˜83 yields
@@
where nearly all terms cancel each other again due to the fact that we can shift the index within the expectation value. By taking the limit $m\rightarrow\infty$ and using the assumptions (ii) and (iii) (Equations˜77 and 78) we have
@@
Finally, by dividing Equation˜86 by Equation˜84 we arrive at the final expression
@@
Appendix D Calculation of the Noise Autocorrelation
We want to calculate the autocorrelation function of epoch-based SGD for a fixed weight vector $\boldsymbol{\mathbf{\theta}}$ and under the assumption that the total number of examples is an integer multiple of the number of examples per batch. For that we repeat the following definitions:
$\boldsymbol{\mathbf{\delta g}}_{k}(\boldsymbol{\mathbf{\theta}})$ $\coloneqq\frac{1}{S}\sum_{n\in\mathcal{B}_{k}}\boldsymbol{\mathbf{\nabla}}\bigl(l(\boldsymbol{\mathbf{\theta}},x_{n})-L(\boldsymbol{\mathbf{\theta}})\bigr)$ (88)
@@ Appendix D Calculation of the Noise Autocorrelation
$N$  $\dots\text{ total number of examples}$ (91)
$S$  $\dots\text{ number of examples per batch}$ (92)
@@
$\boldsymbol{\mathbf{\delta g}}_{k}(\boldsymbol{\mathbf{\theta}})$ $=\frac{1}{S}\sum_{n\in\mathcal{B}_{k}}\boldsymbol{\mathbf{\delta g_{e}}}(n,\boldsymbol{\mathbf{\theta}})$
$=\frac{1}{S}\sum_{n=1}^{N}\boldsymbol{\mathbf{\delta g_{e}}}(n,\boldsymbol{\mathbf{\theta}})s_{k}^{n}$ (93)
@@
$=\begin{cases}1\ \text{ if }~n\in\mathcal{B}_{k}\\ 0\ \text{ if }~n\notin\mathcal{B}_{k}\end{cases}$ (94)
$\boldsymbol{\mathbf{\delta g_{e}}}(n,\boldsymbol{\mathbf{\theta}})$ $\coloneqq\boldsymbol{\mathbf{\nabla}}\bigl(l(\boldsymbol{\mathbf{\theta}},x_{n})-L(\boldsymbol{\mathbf{\theta}})\bigr)\ .$ (95)
@@
$\textrm{cov}\bigl(\vec{\delta g}_{k}(\vec{\theta}),\vec{\delta g}_{k+h}(\vec{\theta})\bigr)$ $=\mathbb{E}\left[\boldsymbol{\mathbf{\delta g}}_{k}(\boldsymbol{\mathbf{\theta}})\boldsymbol{\mathbf{\delta g}}_{k+h}(\boldsymbol{\mathbf{\theta}})^{\top}\right]$
$=\frac{1}{S^{2}}\sum_{n,\tilde{n}=1}^{N}\boldsymbol{\mathbf{\delta g_{e}}}(n,\boldsymbol{\mathbf{\theta}})\boldsymbol{\mathbf{\delta g_{e}}}(\tilde{n},\boldsymbol{\mathbf{\theta}})^{\top}\mathbb{E}\left[s_{k}^{n}s_{k+h}^{\tilde{n}}\right$]\ . (96)
@@
The first term of the right hand side of Equation˜99 is the probability that a given example occurs in a batch, assuming that we already know one of the examples of that batch.
[…]
$=\frac{N-S}{N-1}\,\delta_{n,\tilde{n}}+\textrm{const.}$ (100)
The second term of Equation˜99 is multiplied by $(1-\delta_{h,0})$. Therefore, we assume $h\geq 1$ for the following argument. That is, we want to know the probabilities under the assumption that we are comparing examples from different batches. If the two batches are still from the same epoch, examples cannot repeat as the total number of examples is an integer multiple of the number of examples per batch and because of that every example is shown only once per epoch. Therefore, for $h\geq 1$ holds:
[…]
$=-\frac{S}{N-1}\delta_{n,\tilde{n}}+\textrm{const.}$ (101)
@@
$\mathbb{E}\left[s_{k}^{n}s_{k+h}^{\tilde{n}}\right$] $=\delta_{n,\tilde{n}}\,\frac{S}{N}\frac{N-S}{N-1}\left(\delta_{h,0}-\mathbf{1}_{\{1,\dots,M\}}(h)\frac{S}{N-S}\frac{M-h}{M}\right)+\textrm{const.}$
$=\delta_{n,\tilde{n}}\,S^{2}\left(\frac{1}{S}-\frac{1}{N}\right)\frac{}{}\frac{1}{N-1}\left(\delta_{h,0}-\mathbf{1}_{\{1,\dots,M\}}(h)\frac{M-h}{M\,(M-1)}\right)+\textrm{const.}$ (104)
@@
$\sum_{n,\tilde{n}=1}^{N}\boldsymbol{\mathbf{\delta g_{e}}}(n,\boldsymbol{\mathbf{\theta}})\boldsymbol{\mathbf{\delta g_{e}}}(\tilde{n},\boldsymbol{\mathbf{\theta}})^{\top}\,\delta_{n,\tilde{n}}$ $=\sum_{n=1}^{N}\boldsymbol{\mathbf{\delta g_{e}}}(n,\boldsymbol{\mathbf{\theta}})\boldsymbol{\mathbf{\delta g_{e}}}(n,\boldsymbol{\mathbf{\theta}})^{\top}$
$\eqqcolon(N-1)\,\boldsymbol{\mathbf{C}}_{0}(\boldsymbol{\mathbf{\theta}})\ ,$ (105)
$\sum_{n,\tilde{n}=1}^{N}\boldsymbol{\mathbf{\delta g_{e}}}(n,\boldsymbol{\mathbf{\theta}})\boldsymbol{\mathbf{\delta g_{e}}}(\tilde{n},\boldsymbol{\mathbf{\theta}})^{\top}\cdot 1$ $=\left(\sum_{n=1}^{N}\boldsymbol{\mathbf{\delta g_{e}}}(n,\boldsymbol{\mathbf{\theta}})\right)\left(\sum_{n=1}^{N}\boldsymbol{\mathbf{\delta g_{e}}}(n,\boldsymbol{\mathbf{\theta}})^{\top}\right)$
$=\boldsymbol{\mathbf{\nabla}}\bigl(L(\boldsymbol{\mathbf{\theta}})-L(\boldsymbol{\mathbf{\theta}})\bigr)\boldsymbol{\mathbf{\nabla}}^{\top}\bigl(L(\boldsymbol{\mathbf{\theta}})-L(\boldsymbol{\mathbf{\theta}})\bigr)$
@@
we can insert the expectation value $\mathbb{E}\left[s_{k}^{n}s_{k+h}^{\tilde{n}}\right$] into Equation˜96 and arrive at the final expression: